# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v1

@@ E.1.1 Preliminary results
As the Jacobian of $F$ has full rank almost everywhere, the above set is a null set. Hence, almost all initializations with $Q>8$ diverge. When $Q<8$, we have that the orbit converges to $\left\{z=0\right\}$ or to $\left\{Q=4|\mu|\right\}=\left\{w=2\mathrm{sgn}(\mu)(z+\mu),w\leq 4|\mu|\right\}$. Notice that, when $w=2\mathrm{sgn}(\mu)(z+\mu)$, the $w$-update is given by
$w^{\prime}=\kappa(w)=\frac{1}{4}w\Big(w-2-2|\mu|\Big)^{2}.$
When confined on $w\in[0,4|\mu|]$, the above map has two fixed point $w=0,w=2|\mu|$. It’s easy to verify that, when $|\mu|<1$, $w=0$ is repelling, $w=2|\mu|$ is attracting, and all orbits in this one-dimensional system converges to $w=2|\mu|$ except that with initial value $w=0$. Note $w=0$ corresponds to the fixed point $(-\mu,0)$ of $F$. The Jacobian of $F$ at $(-\mu,0)$ has eigenvalues $(1+\mu)^{2},(-1+\mu)^{2}$. Therefore this point is a hyperbolic fixed point. By the local stable manifold theorem and the fact that $\left\{w=-\mathrm{sgn}(\mu)(z+\mu)\right\}$ is invariant under $F$, we have that the basin of attraction of $(-\mu,0)$ can be given by
$\cup_{N=0}^{\infty}F^{-N}(O\cap\left\{w=-\mathrm{sgn}(\mu)(z+\mu)\right\}),$
for some small neighborhood $O$ of $(-\mu,0)$. Notice that this set is a null set, since the Jacobian of $F$ has full rank almost everywhere. Therefore, for almost all initializations with $Q<8$, we have that they converge to $\left\{z=0\right\}$ or to $\left\{Q=4|\mu|\right\}$, but not to $(-\mu,0)$. Since $w=2|\mu|$ attracts all orbits except $(-\mu,0)$ on the invariant set $\{Q=4|\mu|\}$, we have that all these orbits that converge to $\left\{Q=4|\mu|\right\}$ but not to $(-\mu,0)$ must converge to $(0,2|\mu|)\in\left\{z=0\right\}$. Hence, almost all initializations with $Q<8$ converge to $\left\{z=0\right\}$.
Now consider $|\mu|>1$. We have that $\inf\left\{w\colon(0,w)\in\Omega\right\}=2|\mu|$. Therefore, if $z\neq 0$,
$|z^{\prime}/z|=|z^{2}+\mu z-w+1|\geq|w-1|>1.$
This implies that converging to any global minimizer can only occurs within finite steps, which is a measure-zero event as shown above. This completes the proof. ∎
Note, the above proof implies the following corollary.
@@
Next, we analyze dynamics on the boundary $\{w=\mu z+4\}$. By Proposition 14, the boundary is forward-invariant. Moreover, the system reduces to a one dimensional system
@@
The system $\tilde{F}$ on I=[2,2]I=[-2,2] is Devaney-chaotic and has topological entropy $\log 3$. Moreover, there exists periodic orbits with any period and thus $\tilde{F}$ is also Li-Yorke chaotic.
@@
Similarly, one can verify that $\tilde{F}_{1}\circ\psi=\psi\circ\tilde{F}_{2}$ also holds on (1/3,1/3)(-1/3,1/3) and $[1/3,1]$. Hence, $\tilde{F}_{2}$ is topologically conjugate to $\tilde{F}$, and $\tilde{F}$ is chaotic if and only if $\tilde{F}_{2}$ is (Elaydi,, 2007, Theorem 3.9,).
Note $\tilde{F}_{2}$ is a piecewise linear continuous map with slope equal to $\pm 3$. Hence, the topological entropy of $\tilde{F}_{2}$ is equal to $\log 3$ (De Melo and Van Strien,, 2012, Corollary of Theorem 7.2,). For univariate map on a compact interval, positive topological entropy implies Devaney-chaotic (Theorem 3.13, Elaydi,, 2007). Also, topological conjugacy preserves topological entropy (Theorem 1.7, Ch.8, Robinson,, 1998). Therefore, we have $\tilde{F}$ is Devaney chaotic and has topological entropy $\log 3$.
We now show the existence of periodic orbit with any period. According to the Li-Yorke Theorem (Li and Yorke,, 1975), a sufficient condition is that there exists a point $x$ such that $\tilde{F}_{2}^{3}(x)\leq x<\tilde{F}_{2}(x)<\tilde{F}_{2}^{2}(x)$. Such point can be found explicitly. Consider x=5/7x=-5/7. We have that f(x)=1/7f(x)=-1/7, $f^{2}(x)=3/7$, and $f^{3}(x)=-5/7$. This completes the proof. ∎
The topological entropy of the boundary dynamics provide a lower bound of the original gradient descent dynamics.
@@
The set $\partial\mathcal{D}_{\eta}^{\prime}$ is invariant under $\mathrm{GD}_{\eta}$. We have that
$h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})\geq\log 3.$
@@
By Lemma 12, $\partial\mathcal{D}_{\eta}^{\prime}$ is invariant under $\mathrm{GD}_{\eta}$. Notice that the map $(u,v)\mapsto(z,w)$ is a semi-conjugacy between $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}}$ and $\tilde{F}$. Therefore, with Theorem 1.7, Chapter 8 in Robinson, (1998) we have $h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})\geq h(\tilde{F})=\log 3$. Since $\partial\mathcal{D}_{\eta}^{\prime}$ is an invariant subset, with Proposition 8.1.7 in Vries, (2014), we have that $h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})$, which completes the proof. ∎
@@
Notice that
$u^{\prime}+v^{\prime}=u-\eta zv-\eta\lambda u+v-\eta zu-\eta\lambda v=(1-\eta z-\eta\lambda)(u+v).$
Therefore, $u_{k}+v_{k}$ is always parallel to $u_{0}+v_{0}$. Consider the map $\tau\colon\mathbb{R}^{2d}\to\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$:
@@
As in Proposition 13, let $G_{0}\colon\Omega\to\mathrm{cl}(A_{0})$ denote the inverse of $F|_{\mathrm{cl}(A_{0})}$. Let $m_{0}=(z^{*},w^{*})$ and $m_{k}=G_{0}^{k}(m_{0})$ for $k\geq 1$. We show that $\lim_{k\to\infty}m_{k}=m^{*}=(-2,4-2\mu)$. Note for all $k\geq 1$, $m_{k}\notin\left\{Q=4|\mu|\right\}$. Therefore, by Lemma 12, we know that $Q(m_{k})$ either stays at $8$ or is monotonic. Hence, as we shown in the proof of Proposition 14, $m_{k}$ must converge to
@@
Therefore, $m_{k}$ must converge to the set $Z\cap A_{0}$. Intersecting $Z$ with $\mathrm{cl}(A_{0})$, we get
$Z\cap A_{0}=\left\{z\leq-1,w=\mu z+4\right\}.$
Notice that, when restricted to $\{w=\mu z+4\}$, the system $F$ reduces to $\tilde{F}(z)=z^{3}-3z,\quad z\in[-2,2]$. Then $G_{0}|_{Z\cap A_{0}}$ corresponds to the branch of $\tilde{F}^{-1}$ whose image is [2,1][-2,-1]. Using the conjugacy we provided in Proposition 17, it’s easy to obtain that z=2z=-2 is the unique, globally attracting fixed point for the map $G_{0}|_{Z\cap A_{0}}$. Since $m_{k}$ converges to $Z\cap A_{0}$ and z=2z=-2 is globally attracting, we have that $m_{k}\to m^{*}$ as $k\to\infty$.
Now consider the given open set $O$ which satisfies $O\cap\left\{Q=8\right\}$ is not empty. We prove that there exists $x\in O$ and $n$ such that $F^{n}(x)=m^{*}$. Notice that $F|_{\left\{Q=8\right\}}$ is topologically conjugate to the piece-wise linear map $\tilde{F}_{2}$ as we shown in the proof of Proposition 17. For $\tilde{F}_{2}$, the full preimage of the two endpoint of the interval [1,1][-1,1] is given by
$\cup_{k\geq 0}\{-1+\frac{2j}{3^{k}}\colon j=0,1,\cdots,3^{k}\},$
which is dense in [1,1][-1,1]. Particularly, it is clear that the preimage of 1-1 is dense. Therefore, there exists $x\in O$ and $n$ such that $F^{n}(x)=m^{*}$, i.e., $x\in F^{-n}(m^{*})$. Notice $\left\{Q=8\right\}\subset\mathrm{cl}(B)$, by Proposition 13, we have that there exists $i_{1},\cdots,i_{n_{1}}\in\left\{0,1,2\right\}$ such that $G_{i_{n_{1}}}\circ\cdots\circ G_{i_{1}}(m^{*})=x.$ Since the composition of $G_{i}$’s is continuous, there exists a neighborhood $\tilde{O}$ of $m^{*}$ such that $G_{i_{n_{1}}}\circ\cdots\circ G_{i_{1}}(\tilde{O})\subset O$. Since $m_{k}\to m^{*}$, there exists $n_{2}$ such that $m_{n_{2}}=G_{0}^{n_{2}}(0,w^{*})\in\tilde{O}$. Taken together, we have that
$G_{i_{n_{1}}}\circ\cdots\circ G_{i_{1}}\circ G_{0}^{n_{2}}(0,w^{*})\triangleq\hat{m}\in O.$
This implies that
$F^{n_{1}+n_{2}}(\hat{m})=(0,m^{*}),$

### v2

@@ E.1.1 Preliminary results
As the Jacobian of $F$ has full rank almost everywhere, the above set is a measure-zero set (Ponomarev,, 1987). Hence, almost all initializations with $Q>8$ diverge.
When $Q<8$, by Proposition 14, we have that the orbit converges to $\left\{z=0\right\}$ or to $\left\{Q=4|\mu|\right\}=\left\{w=2\mathrm{sgn}(\mu)(z+\mu),w\leq 4|\mu|\right\}$. Notice that, when $w=2\mathrm{sgn}(\mu)(z+\mu)$, the $w$-update under $F$ is given by
$w^{\prime}=\kappa(w)=\frac{1}{4}w\Big(w-2-2|\mu|\Big)^{2}.$ (19)
Consider $\kappa$ as a one-dimensional dynamical system defined on $0,4|\mu|]$. For this one-dimensional system, it is straight forward to obtain that, there are two fixed points: $w=0$ and $w=2|\mu|$, and also that, when $|\mu|<1$, all orbits converge to $w=2|\mu|$ except the one with initial value $w=0$. Note that $w=0$ corresponds to the fixed point $(-\mu,0)$ of $F$. The Jacobian of $F$ at $(-\mu,0)$ has eigenvalues $(1+\mu)^{2},(-1+\mu)^{2}$. Therefore, $(-\mu,0)$ is a hyperbolic fixed point. By the local stable manifold theorem and the fact that $\left\{w=-\mathrm{sgn}(\mu)(z+\mu)\right\}$ is invariant under $F$, we have that the basin of attraction of $(-\mu,0)$ can be given by
$B(-\mu,0)=\cup_{N=0}^{\infty}F^{-N}(O\cap\left\{w=-\mathrm{sgn}(\mu)(z+\mu)\right\}),$
for some small neighborhood $O$ of $(-\mu,0)$. This set is a measure-zero set, since the Jacobian of $F$ has full rank almost everywhere. Therefore, for all initializations lies in $\left\{Q(z,w)<8\right\}\setminus B(-\mu,0)$, the orbit converges to $\left\{z=0\right\}$ or, to $\left\{Q=4|\mu|\right\}\setminus\{(-\mu,0)\}$. Consider any fixed initialization $(z_{0},w_{0})$ in the second case. Since $(-\mu,0)$ is a hyperbolic fixed point, the orbit does not have an accumulation point in some neighborhood of $(-\mu,0)$. Therefore, the omega-limit set $\omega(z_{0},w_{0})\subset\left\{Q=4|\mu|\right\}\setminus\{(-\mu,0)\}$. The omega-limit set is non-empty, as the orbit is always bounded. Note, for any $m\in\omega(z_{0},w_{0})$, we have that $F^{N}(m)\to(0,-2|\mu|)$ as $N$ tends to infinity, as explained above. Since the omega-limit set is forward invariant under $F$ and closed, we have that $(0,-2|\mu|)\in\omega(z_{0},w_{0})$. It follows that the orbit visits an arbitrarily small neighborhood $O$ of $(0,-2|\mu|)$ at some time $k$. Take the neighborhood as $O=\left\{|Q(z,w)-4|\mu||<\varepsilon,|z|<\varepsilon\right\}$. Notice that
$|\frac{z_{k+1}}{z_{k}}|=|z_{k}^{2}+\mu z_{k}-w_{k}+1|\leq|z_{k}^{2}+\mu z_{k}|+|w_{k}-1|.$
Since $|z^{2}+\mu z|+|w-1|\to|2|\mu|-1|$ as $(z,w)\to(0,2|\mu|)$, and $|2|\mu|-1|<1$, we can choose $\varepsilon$ small enough such that $|\frac{z_{k+1}}{z_{k}}|<1-\delta$ for some $\delta>0$. Therefore, $|z_{k+1}|<|z_{k}|$. Meanwhile, note that the $Q$ value monotonically decreases as the orbit is in $\{Q<8\}$. Hence, for all $j\geq k$, $(z_{j},w_{j})\in O$ and $|\frac{z_{j+1}}{z_{j}}|<1-\delta$. It follows that $z_{j}\to 0$ and the orbit converges to $\left\{z=0\right\}$. Therefore, we have that almost all initializations with $Q<8$ converge to $\left\{z=0\right\}$.
Finally, consider $|\mu|>1$. Since $2|\mu|=2\mathrm{sgn}(\mu)(0+\mu)$, we have that $\inf\left\{w\colon(0,w)\in\Omega\right\}=2|\mu|>2$. Using arguments similar to those in Proposition 14, in particular in (18), we have that converging to any global minimizer must occur within finite steps. As shown above, those initializations form a measure-zero set. This completes the proof. ∎
From the proofs of the preceding two propositions, we obtain the following corollary.
@@
Next, we analyze dynamics on the invariant set $\left\{Q(z,w)=8\right\}=\{w=\mu z+4\}$. Observe that when restricted to this set, $F$ reduces to the following one-dimensional system:
@@
The following result shows that $\tilde{F}$ is a chaotic dynamical system.
@@
Assume $|\mu|<1$. The system $\tilde{F}$ on I=[2,2]I=[-2,2] is Devaney chaotic and has topological entropy $\log 3$. Moreover, there exists periodic orbits with any period and thus $\tilde{F}$ is also Li-Yorke chaotic.
@@
Similarly, one can verify that $\tilde{F}_{1}\circ\psi=\psi\circ\tilde{F}_{2}$ also holds on (1/3,1/3)(-1/3,1/3) and $[1/3,1]$. Hence, $\tilde{F}_{2}$ is topologically conjugate to $\tilde{F}$.
Note $\tilde{F}_{2}$ is a piecewise linear continuous map with slope equal to $\pm 3$. Hence, the topological entropy of $\tilde{F}_{2}$ is equal to $\log 3$ (De Melo and Van Strien,, 2012, Corollary of Theorem 7.2,). For a univariate map on a compact interval, positive topological entropy implies Devaney chaotic (Theorem 3.13, Elaydi,, 2007). Due to the conjugacy, we have that $\tilde{F}$ is also Devaney chaotic. Also, conjugacy preserves topological entropy (Theorem 1.7, Ch.8, Robinson,, 1998). Therefore, $h(\tilde{F})=\log 3$.
We now show the existence of periodic orbit with any period. According to the Li-Yorke Theorem (Li and Yorke,, 1975), a sufficient condition is that there exists a point $x$ such that $\tilde{F}_{2}^{3}(x)\leq x<\tilde{F}_{2}(x)<\tilde{F}_{2}^{2}(x)$. Consider x=5/7x=-5/7. We have that $\tilde{F}_{2}(x)=-1/7$, $\tilde{F}_{2}^{2}(x)=3/7$, and $\tilde{F}_{2}^{3}(x)=-5/7$. Due to the conjugacy, $\tilde{F}$ also has periodic orbit with any period and is Li-Yorke chaotic. This completes the proof. ∎
We translate Proposition 17 to the original gradient system $\mathrm{GD}_{\eta}$.
@@
Assume $|\mu|<1$. We have that $h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})\geq\log 3$, and $\mathrm{GD}_{\eta}$ admits periodic orbits with any period.
@@
By Lemma 12, $\partial\mathcal{D}_{\eta}^{\prime}$ is invariant under $\mathrm{GD}_{\eta}$. Since the map $T(u,v)=(z,w)$ is a semi-conjugacy between $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}}$ and $\tilde{F}$, we have that $h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})\geq h(\tilde{F})=\log 3$ (see, e.g., Theorem 1.7, Ch 8. Robinson,, 1998), where the equality comes from Proposition 17. Meanwhile, since $\partial\mathcal{D}_{\eta}^{\prime}$ is an invariant subset of $\mathbb{R}^{2d}$, we have that $h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})$ (see, e.g., Proposition 8.1.7. Vries,, 2014).
Next, we show the existence of all periodic orbits. As shown in Proposition 17, $\tilde{F}$ is conjugate to $\tilde{F}_{2}$:
$\tilde{F}_{2}(z)=\begin{cases}3z+2,\quad&x\in I_{0}=[-1,-1/3];\\ -3z,\quad&x\in I_{1}=(-1/3,1/3);\\ 3z-2,\quad&x\in I_{2}=[1/3,1].\end{cases}$ (20)
Note $\tilde{F}_{2}$ is a piecewise linear map defined on [1,1][-1,1], with each piece mapping onto the whole interval [1,1][-1,1]. With classical analyses in symbolic dynamics (see, e.g., Ch 1.7. Devaney and Eckmann,, 1987), we have that $\tilde{F}_{2}$, when restricted to $1,1]-\{x\in[-1,1]\colon\tilde{F}_{2}^{N}(x)\neq\pm 1,\forall N\}$, is conjugate to the shift operator $\sigma$, defined on the following set
[…]
such that $\sigma(\boldsymbol{s})=\sigma(s_{0}s_{1}s_{2}\cdots)=(s_{1}s_{2}\cdots)$. The conjugacy is given by the map $\zeta$:
$\zeta(x)=(s_{0}s_{1}\cdots),\quad s_{j}=l\ \text{ if }\tilde{F}_{2}^{j}(x)\in I_{l},$
where $l\in\left\{0,1,2\right\}$ and $I_{l}$ is defined in (20).
Now we construct periodic orbits $(u_{k},v_{k})_{k\geq 0}$ with periodicity $K$ for arbitrary $K$. Consider $T(u_{k},v_{k})=(z_{k},w_{k})$. Consider new coordinates in $uv$-space, given by (u+v,uv)(u+v,u-v), which preserves all dynamical behaviors of $\mathrm{GD}_{\eta}$. Notice that
$u_{k+1}+v_{k+1}=u_{k}-\eta z_{k}v_{k}+v_{k}-\eta z_{k}u_{k}=(1-\eta z_{k})(u_{k}+v_{k}),$ (21)
$u_{k+1}-v_{k+1}=u_{k}-\eta z_{k}v_{k}-v_{k}+\eta z_{k}u_{k}=(1+\eta z_{k})(u_{k}-v_{k}).$
It follows that, during training $u_{k}+v_{k}$ and $u_{k}-v_{k}$ must lie in the one-dimensional space spanned by $u_{0}+v_{0}$ and $u_{0}-v_{0}$, respectively. Meanwhile, as shown in Proposition 7, if $T(u^{\prime},v^{\prime})=T(u,v)$, then $\|u^{\prime}+v^{\prime}\|=\|u+v\|$ and $\|u^{\prime}-v^{\prime}\|=\|u-v\|$. Also, it is clear that if $T(u^{\prime},v^{\prime})\neq T(u,v)$, $(u^{\prime},v^{\prime})\neq(u,v)$. Therefore, whenever $(z_{k},w_{k})_{k\geq 0}$ is a $K$-orbit and,
$\left\{k\in\left\{0,\cdots,K-1\right\}\colon 1-\eta z_{k}<0\right\},\quad\#\left\{k\in\left\{0,\cdots,K-1\right\}\colon 1+\eta z_{k}<0\right\}$
are two even numbers, where #S\#S is the number of elements in the set $S$, we have that $(u_{k},v_{k})_{k\geq 0}$ is a $K$-orbit. Notice that, under the conjugacy between $\tilde{F}$ and $\tilde{F}_{2}$, these two numbers correspond to the numbers of visits of the $\tilde{F}_{2}$-orbit to $I_{2}$ and $I_{0}$, i.e., the number of $2$’s and 0’s appearing in the symbol $\boldsymbol{s}$. For $K=2$, take $(z_{0},w_{0})=(-2,4-2\mu)$. Notice that $(-2,4-2\mu)$ is a fixed point under $F$, and that, $1-\eta z_{0}-\eta\lambda>0$ but $1+\eta z_{0}-\eta\lambda<0$. Therefore, $(u_{k},v_{k})_{k\geq 0}$ is a $2$-orbit. For $K=3$, take $(z_{0},w_{0})$ be the element corresponding to (001 001)($001\cdots$), which is a $3$-orbit in the symbolic system. As there are two 0’s and zero $2$’s, $(u_{k},v_{k})_{k\geq 0}$ is a $3$-orbit. For $K=4$, take $(z_{0},w_{0})$ be the element corresponding to (0011 0011)($0011\cdots$), and we have that $(u_{k},v_{k})_{k\geq 0}$ is a $4$-orbit. For $K=2N+1$, take $(z_{0},w_{0})$ corresponding to the repetition of $2N$ 0’s and one $1$; for $K=2N+2$, take $(z_{0},w_{0})$ corresponding to the repetition of $2N$ 0’s and two $1$’s. By this, we have that $\mathrm{GD}_{\eta}$ admits all periodic orbits, which completes the proof. ∎
@@
Let $(u_{k},v_{k})_{k\geq 0}$ be an orbit under $\mathrm{GD}_{\eta}$. As shown in (21), $u_{k}+v_{k}$ is always parallel to $u_{0}+v_{0}$. Consider the map $\tau\colon\mathbb{R}^{2d}\to\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$:
@@
As in Proposition 13, let $G_{0}\colon\Omega\to\mathrm{cl}(A_{0})$ denote the inverse of $F|_{\mathrm{cl}(A_{0})}$. Let $m_{0}=(z^{*},w^{*})$ and $m_{k}=G_{0}^{k}(m_{0})$ for $k\geq 1$. We first show that $\lim_{k\to\infty}m_{k}=m^{*}=(-2,4-2\mu)$. Note for all $k\geq 1$, $m_{k}\notin\left\{Q=4|\mu|\right\}$. Therefore, by Lemma 12, we know that $Q(m_{k})$ either stays at $8$ or monotonically approaches $8$. Hence, as we shown in the proof of Proposition 14, $m_{k}$ must converge to the compact set
@@
As $G_{0}(\Omega)\subset\mathrm{cl}(A_{0})$, $m_{k}$ must converge to the set
$Z\cap\mathrm{cl}(A_{0})=\left\{z\leq-1,w=\mu z+4\right\}.$
It follows that the omega-limit set $\omega(m_{0})\subset Z\cap\mathrm{cl}(A_{0})$. Notice that, when restricted to $\{w=\mu z+4\}$, the system $F$ reduces to $\tilde{F}(z)=z^{3}-3z,\quad z\in[-2,2]$. Then $G_{0}|_{Z\cap\mathrm{cl}(A_{0})}$ corresponds to the branch of $\tilde{F}^{-1}$ whose image is [2,1][-2,-1]. Using the conjugacy between $\tilde{F}$ and $\tilde{F}_{2}$ as shown in Proposition 17, it’s clear that all orbits under $G_{0}|_{Z\cap\mathrm{cl}(A_{0})}$ converge to z=2z=-2. Therefore, for any $q\in\omega(m_{0})$, $G_{0}^{N}(q)\to(-2,4-2\mu)$ as $N\to\infty$. Since the omega-limit set is invariant and closed, we have that $(-2,4-2\mu)\in\omega(m_{0})$, which implies that $m_{k}$ visits an arbitrarily small neighborhood of $(-2,4-2\mu)$. Note, the eigenvalues of $JF(-2,4-2\mu)$ are $9$ and $5-2\mu$. Hence, $(-2,4-2\mu)$ attracts all orbits of $G_{0}$ in some neighborhood of itself. Hence, $m_{k}\to m^{*}$ as $k\to\infty$.
Now consider any open set $O$ that satisfies $O\cap\left\{Q=8\right\}=O\cap\left\{w=\mu z+4\right\}$ is not empty. We show that there exists $x\in O$ and $n$ such that $F^{n}(x)=m^{*}$. Recall that $F|_{\left\{Q=8\right\}}$ is topologically conjugate to the piece-wise linear map $\tilde{F}_{2}$ as shown in Proposition 17. For $\tilde{F}_{2}$, we have that
$\cup_{N=0}^{\infty}\tilde{F}_{2}^{-N}(\left\{\pm 1\right\})=\cup_{k\geq 0}\{-1+\frac{2j}{3^{k}}\colon j=0,1,\cdots,3^{k}\},$
which is dense in [1,1][-1,1]. By symmetry arguments, we have that the preimage of 1-1 is dense. Therefore, using the conjugacy, we have that there exists $x\in O$ and $n$ such that $x\in F^{-n}(m^{*})$, i.e., $F^{n}(x)=m^{*}$. Notice $\left\{Q=8\right\}\subset\mathrm{cl}(B)$, by Proposition 13, we have that there exists $i_{1},\cdots,i_{n_{1}}\in\left\{0,1,2\right\}$ such that $G_{i_{n_{1}}}\circ\cdots\circ G_{i_{1}}(m^{*})=x.$ Since the composition of $G_{i}$’s is continuous, there exists a neighborhood $\tilde{O}$ of $m^{*}$ such that $G_{i_{n_{1}}}\circ\cdots\circ G_{i_{1}}(\tilde{O})\subset O$. Since $m_{k}\to m^{*}$, there exists $n_{2}$ such that $m_{n_{2}}=G_{0}^{n_{2}}(m_{0})\in\tilde{O}$. Taken together, we have that
$G_{i_{n_{1}}}\circ\cdots\circ G_{i_{1}}\circ G_{0}^{n_{2}}(m_{0})\triangleq\hat{m}\in O.$
By the definitions of $G_{i}$’s as in Proposition 13, it follows that
$F^{n_{1}+n_{2}}(\hat{m})=m_{0},$